# Characterizing graph-nonedge pairs with single interval Cayley configuration spaces in 3-dimensions

### v1

@@ C.4. Proof of Proposition 9 [Pruned possible graphs for top-level $H^{-1}$]
In the case where $f^{\prime}=w_{1}w_{3}$, we first argue that some atom of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$ contains $w_{1}w_{4}$ and has a $w_{1}w_{4}$-preserving forbidden minor. By Lemma 50, $J$ contains $w_{4}$. Hence, since $J$ is an atom, each cms and atom of $J\cup w_{1}w_{4}$ contains $w_{1}w_{4}$. Combining this with the fact that $J\cup w_{1}w_{4}$ has a $w_{1}w_{3}$-preserving forbidden minor and Lemma 27 shows that some atom of $J\cup w_{1}w_{4}$, and therefore some atom $K$ of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$, contains $w_{1}w_{4}$ and has a $w_{1}w_{3}$-preserving forbidden minor $[K]_{A}$. If $w_{1}w_{4}$ is preserved in $[K]_{A}$, then the claim is true. Else, we define a vertex $x$ as follows. If $K$ contains $w_{3}$, then set $x=w_{3}$. Otherwise, some cms $E$ of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$ contained in $K$ is such that all paths between $w_{3}$ and any vertex in $K\setminus E$ contains some vertex in $E$. Since $[K]_{A}$ preserves $w_{1}w_{3}$ but not $w_{1}w_{4}$, $E\setminus\{w_{1},w_{4}\}$ contains some vertex $x$ such that $[x]_{A}=[w_{3}]_{A}$.
@@
Next, for either case of $x$, there clearly exists a $w_{1}x$-preserving minor $[K]_{B}$ from which $[K]_{A}$ can be obtained by contracting $[w_{1}w_{4}]_{B}$. Hence, $[K]_{B}$ is one of the graphs in Figures 8, 17, or 18 such that $[w_{1}w_{4}]_{B}$ is the green edge. Since $K$ is an atom, it is $2$-connected, and so Lemma 10 allows us to assume that $[K]_{B}$ is one of the graphs in Figures 8 or 18. Furthermore, since $[K]_{A}$ preserves $w_{1}w_{3}$ but not $w_{1}w_{4}$, $[w_{3}]_{B}$ is neither $[w_{1}]_{B}$ nor $[w_{4}]_{B}$. Combining this with the fact that $w_{3}$ neighbors both $w_{1}$ and $w_{4}$ shows that $[K]_{B}$ cannot be any graph in Figure 8. Therefore, Lemma 11 shows that $[K]_{B}$ has a $[w_{1}w_{4}]_{B}$-preserving forbidden minor, and hence $K$ has a $w_{1}w_{4}$-preserving forbidden minor.
@@
Finally, we show that there exists some atom of $[G\cup f]$ contains $[f]$ and has an $[f]$-preserving forbidden minor. Note that $K\setminus w_{1}w_{3}$ is some induced subgraph $N$ of $[G^{\prime}]$ such that $w_{1}w_{4}$ is $[w_{1}w_{4}]$. If $K$ does not contain $w_{3}$, then $N$ is an atom that contains $[f]$ and has an $[f]$-preserving forbidden minor, since $[f]=[w_{1}w_{4}]$. Hence, the atom of $[G\cup f]$ that contains $N$ has the desired properties. Otherwise, if $K$ contains $w_{3}$, then let $G_{1}=[H^{-1}]$, $G_{2}=N$, $S=G_{1}\cap G_{2}$, and $e=[w_{1}w_{3}]$. Observe that $G_{1}\cup G_{2}$ contains $[f]$ and has an $[f]$-preserving forbidden minor. Lastly, since $G_{1}$ and $G_{2}\cup e$ are atoms, Lemma 64 shows that $G_{1}\cup G_{2}$ is an atom. Therefore, the atom of $[G\cup f]$ that contains $G_{1}\cup G_{2}$ has the desired properties. ∎
Lemma 53 ($H^{-1}$ is not graph in Figures 15(g) or 15(h)).
@@
Neither of the graphs in Figures 15(g) or 15(h) such that $uw_{1}$ is an edge is the expanded $f$-component of any top-level expanded $f$-separating cms of a minimal graph-nonedge pair.
@@
Assume to the contrary that one of the graphs in Figures 15(g) or 15(h) such that $uw_{1}$ is an edge is the expanded $f$-component $H^{-1}$ of some top-level expanded $f$-separating cms $C^{-1}$ of a minimal graph-nonedge pair $(G,f)$. We will show that either $uw_{1}$ or $w_{2}w_{4}$ is reducing with respect to $(G,f)$, which contradicts the fact that $(G,f)$ is minimal. Observe that $f$ is retained in the minor $[G\cup f]$ obtained by contracting $uw_{1}$, and so $uw_{1}$ is not of Type (3). We attempt to show that $uw_{1}$ is not of Type (1) or (2) by identifying some atom of $[G\cup f]$ contains $[f]$ and has an $[f]$-preserving forbidden minor. Let $G^{\prime}=G\setminus(H^{-1}\setminus C^{-1})$. It suffices to show that some atom of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$ contains $w_{1}w_{4}$ and has a $w_{1}w_{4}$-preserving forbidden minor. By Lemma 47, there exists a nonedge $f^{\prime}$ of $C^{-1}$ such that some atom $J$ of $G^{\prime}\cup f^{\prime}$ contains $f^{\prime}$ and has an $f^{\prime}$-preserving forbidden minor and the pair $(G^{\prime},f^{\prime})$ has no reducing edge. If $f=w_{1}w_{3}$, then the argument is identical to the proof of Lemma 52, so we omit the details.
@@
Next, consider the case where $f^{\prime}=w_{1}w_{4}$. By Lemma 50, $J$ contains $w_{3}$. Hence, since $J$ is an atom, each cms and atom of $J\cup w_{1}w_{3}$ contains $w_{1}w_{3}$. Combining this with the fact that $J\cup w_{1}w_{3}$ has a $w_{1}w_{4}$-preserving forbidden minor and Lemma 27 shows that some atom of $J\cup w_{1}w_{3}$, and therefore some atom $K$ of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$, contains $w_{1}w_{3}$ and has a $w_{1}w_{4}$-preserving forbidden minor $[K]_{A}$. If $K$ contains $w_{4}$, then the claim is true. Otherwise, we show that $w_{2}w_{4}$ is reducing with respect to $(G,f)$. First, we argue that $K$ has a $w_{1}w_{3}$-preserving forbidden minor. If $[K]_{A}$ preserves $w_{1}w_{3}$, then we are done. Else, since $K$ does not contain $w_{4}$, some cms $E$ of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$ contained in $K$ is such that all paths between $w_{4}$ and any vertex in $K\setminus E$ contains some vertex in $E$. Also, since $[K]_{A}$ preserves $w_{1}w_{4}$ but not $w_{1}w_{3}$, $E\setminus\{w_{1},w_{3}\}$ contains some vertex $x$ such that $[x]_{A}=[w_{4}]_{A}$. Therefore, there clearly exists a $w_{1}x$-preserving forbidden minor $[K]_{B}$ from which $[K]_{A}$ can be obtained by contracting $[w_{1}w_{3}]_{B}$. The remainder of the proof that $K$ has a $w_{1}w_{3}$-preserving forbidden minor is identical to part of the proof of Lemma 52, so we omit the details.
@@
Finally, observe that $f$ is retained in the minor $[G\cup f]_{D}$ obtained by contracting $w_{2}w_{4}$, and so $w_{2}w_{4}$ is not of Type (3). We complete the proof by showing that $w_{2}w_{4}$ is not of Type (1) or (2). Since $K$ does not contain $w_{4}$, observe that $K$ is an atom of $G\cup w_{1}w_{3}$ and is weakly retained in the minor of $G\cup w_{1}w_{3}$ obtained by contracting $w_{2}w_{4}$. Hence, $K\setminus w_{1}w_{3}$ is some induced subgraph $N$ of $[G^{\prime}]_{D}$ such that each vertex $w_{i}$ other than $w_{4}$ is $[w_{i}]_{D}$. Let $G_{1}$ be the subgraph of $[G\cup f]_{D}$ induced by $\{[u]_{D},[w_{1}]_{D},[w_{2}]_{D},[w_{3}]_{D}\}$, $G_{2}=N$, $S=G_{1}\cap G_{2}$, and $e=[w_{1}w_{3}]_{D}$. By construction, $G_{1}\cup G_{2}$ contains $[f]_{D}$. Also, it is easy to see that $G_{1}\cup G_{2}$ has an $[f]_{D}$-preserving forbidden minor. We will apply Lemma 65 to show that $G_{1}\cup G_{2}$ is an atom. Note that $e$ is a nonedge of $G_{1}\cup G_{2}$ and both $G_{1}\cup e$ and $G_{2}\cup e$ are atoms. Furthermore, $G_{1}\setminus A$ is connected for any clique subgraph $A$ of $S$. Lastly, the only clique subgraph of $S$ that could separate $[w_{1}]_{D}$ and $[w_{3}]_{D}$ in $G_{2}$ is $[w_{2}]_{D}$. Since $G_{2}\cup e$ has a forbidden minor and $S$ contains at most three vertices, there exists some vertex in $G_{2}\setminus S$. Additionally, since $G_{2}\cup e$ is an atom, there exist paths that connect this vertex to $[w_{1}]_{D}$ and $[w_{3}]_{D}$ that each contain exactly one vertex in $S$. Therefore, $[w_{2}]_{D}$ does not separate $[w_{1}]_{D}$ and $[w_{3}]_{D}$ in $G_{2}$, and so Lemma 65 shows that $K$ is an atom. Thus, $G_{1}\cup G_{2}$ is contained in some atom of $[G\cup f]_{D}$ that contains $[f]_{D}$ and has an $[f]_{D}$-preserving forbidden minor, and the proof is complete. ∎
@@
First, we show that some atom of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$ contains $w_{1}w_{3}$ and has a $w_{1}w_{3}$-preserving forbidden minor. By Lemma 50, $J$ contains $w_{4}$. Hence, since $J$ is an atom, each cms and atom of $J\cup w_{1}w_{4}$ contains $w_{1}w_{4}$. Combining this with the fact that $J\cup w_{1}w_{4}$ has a $w_{1}w_{3}$-preserving forbidden minor and Lemma 27 shows that some atom of $J\cup w_{1}w_{4}$, and therefore some atom $K$ of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$, contains $w_{1}w_{4}$ and has a $w_{1}w_{3}$-preserving forbidden minor $[K]_{A}$. If $K$ contains $w_{3}$, then we are done. Otherwise, we show that $w_{2}w_{3}$ is reducing with respect to $(G,f)$.
Observe that $f$ is retained in the minor $[G\cup f]_{A}$ obtained by contracting $w_{2}w_{3}$, and so $w_{2}w_{3}$ is not of Type (3). To see that $w_{2}w_{3}$ is not of Type (1) or (2), note that an argument similar to to part of the proof of Lemma 53 shows that $K$ has a $w_{1}w_{4}$-preserving forbidden minor. Also, since $K$ does not contain $w_{3}$, $K$ is an atom of $G\cup w_{1}w_{4}$ and is weakly retained in the minor of $G\cup w_{1}w_{4}$ obtained by contracting $w_{2}w_{3}$. Hence, $K\setminus w_{1}w_{4}$ is some induced subgraph $N$ of $[G^{\prime}]_{A}$ such that each vertex $w_{i}$ other than $w_{3}$ is $[w_{i}]_{A}$. Let $G_{1}=[H^{-1}_{A}$, $G_{2}=N$, $S=G_{1}\cap G_{2}$, and $e=[w_{1}w_{4}]_{A}$. By construction, $G_{1}\cup G_{2}$ contains $[f]_{A}$. Furthermore, it is easy to see that $G_{1}\cup G_{2}$ has an $[f]_{A}$-preserving forbidden minor. Lastly, since $e$ is a nonedge of $G_{1}\cup G_{2}$ and both $G_{1}$ and $G_{2}\cup e$ are atoms, Lemma 64 shows that $G_{1}\cup G_{2}$ is an atom. Therefore, $G_{1}\cup G_{2}$ is contained in some atom of $[G\cup f]_{A}$ that contains $[f]_{A}$ and has an $[f]_{A}$-preserving forbidden minor, and the proof is complete.
Finally, since $K$ contains $w_{3}$, $K\setminus w_{1}w_{3}$ is some induced subgraph $N$ of $[G\cup f]$ such that each vertex $w_{i}$ is $[w_{i}]$. Let $G_{1}=[C^{-1}]\cup[u]$, $G_{2}=N$, $S=G_{1}\cap G_{2}$, and $e=[w_{1}w_{3}]$. Observe that $G_{1}\cup G_{2}$ contains $[f]$ and has an $[f]$-preserving forbidden minor. Lastly, since $G_{1}$ and $G_{2}\cup e$ are atoms, Lemma 64 shows that $G_{1}\cup G_{2}$ is an atom. Thus, $G_{1}\cup G_{2}$ is contained in some atom of $[G\cup f]$ that contains $[f]$ and has an $[f]$-preserving forbidden minor, as desired.
@@
By Lemma 50, $J$ contains $w_{2}$ and $w_{3}$. Hence, $J$ is some induced subgraph $G_{2}$ of $[G^{\prime}\cup w_{1}w_{4}]$ such that each vertex $w_{i}$ is $[w_{i}]$. Let $G_{1}=[C^{-1}]\cup[u]$, $S=G_{1}\cap G_{2}$, and $e=[w_{1}w_{3}]$. Observe that $G_{1}\cup G_{2}$ contains $[f]$ and has an $[f]$-preserving forbidden minor. Also, since $e$ is a nonedge and both $G_{1}$ and $G_{2}$ are atoms, Lemma 64 shows that $G_{1}\cup G_{2}$ is an atom. Therefore, $G_{1}\cup G_{2}$ is contained in an atom of $[G\cup f]$ that contains $[f]$ and has an $[f]$-preserving forbidden minor. Thus, $xw_{1}$ is reducing, as desired.

### v2

@@ B.3   Proof of Lemmas for Proposition 9
In the case where $f^{\prime}=w_{1}w_{3}$, we first argue that some atom of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$ contains $w_{1}w_{4}$ and has a $w_{1}w_{4}$-preserving forbidden minor. By Lemma 56, $J$ contains $w_{4}$. Hence, since $J$ is an atom, each CMS and atom of $J\cup w_{1}w_{4}$ contains $w_{1}w_{4}$. Combining this with the fact that $J\cup w_{1}w_{4}$ has a $w_{1}w_{3}$-preserving forbidden minor and Lemma 22 shows that some atom of $J\cup w_{1}w_{4}$, and therefore some atom $K$ of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$, contains $w_{1}w_{4}$ and has a $w_{1}w_{3}$-preserving forbidden minor $[K]_{1}$. If $w_{1}w_{4}$ is preserved in $[K]_{1}$, then the claim is true. Else, we define a vertex $x$ as follows. If $K$ contains $w_{3}$, then set $x=w_{3}$. Otherwise, some CMS $E$ of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$ contained in $K$ is such that all paths between $w_{3}$ and any vertex in $K\setminus E$ contains some vertex in $E$. Since $[K]_{1}$ preserves $w_{1}w_{3}$ but not $w_{1}w_{4}$, $E\setminus\{w_{1},w_{4}\}$ contains some vertex $x$ such that $[x]_{1}=[w_{3}]_{1}$.
@@
Next, for either case of $x$, there clearly exists a $w_{1}x$-preserving minor $[K]_{2}$ from which $[K]_{1}$ can be obtained by contracting $[w_{1}w_{4}]_{2}$. Hence, $[K]_{2}$ is one of the graphs in Figures 8, 14, or 15 with $[w_{1}w_{4}]_{2}$ as the green edge. Since $K$ is an atom, it is $2$-connected, and so Lemma 4 allows us to assume that $[K]_{2}$ is one of the graphs in Figures 8 or 15. Furthermore, since $[K]_{1}$ preserves $w_{1}w_{3}$ but not $w_{1}w_{4}$, $[w_{3}]_{2}$ is neither $[w_{1}]_{2}$ nor $[w_{4}]_{2}$. Combining this with the fact that $w_{3}$ neighbors both $w_{1}$ and $w_{4}$ shows that $[K]_{2}$ cannot be any graph in Figure 8. Therefore, Lemma 5 shows that $[K]_{2}$ has a $[w_{1}w_{4}]_{2}$-preserving forbidden minor, and hence $K$ has a $w_{1}w_{4}$-preserving forbidden minor.
@@
Finally, we show that there exists some atom of $[G\cup f]$ contains $[f]$ and has an $[f]$-preserving forbidden minor. Note that $K\setminus w_{1}w_{3}$ is some induced subgraph $N$ of $[G^{\prime}]$ with $w_{1}w_{4}$ as $[w_{1}w_{4}]$. If $K$ does not contain $w_{3}$, then $N$ is an atom that contains $[f]$ and has an $[f]$-preserving forbidden minor, since $[f]=[w_{1}w_{4}]$. Hence, the atom of $[G\cup f]$ that contains $N$ has the desired properties. Otherwise, if $K$ contains $w_{3}$, then let $G_{1}=[H^{-1}]$, $G_{2}=N$, $S=G_{1}\cap G_{2}$, and $e=[w_{1}w_{3}]$. Observe that $G_{1}\cup G_{2}$ contains $[f]$ and has an $[f]$-preserving forbidden minor. Lastly, since $G_{1}$ and $G_{2}\cup e$ are atoms, Lemma 66 shows that $G_{1}\cup G_{2}$ is an atom. Therefore, the atom of $[G\cup f]$ that contains $G_{1}\cup G_{2}$ has the desired properties. ∎
Lemma 58 ($H^{-1}$ is not graph in Figures 20(g) or 20(h)).
@@
Neither of the graphs in Figures 20(g) or 20(h) such that $uw_{1}$ is an edge is the expanded $f$-component of any top-level expanded $f$-separating CMS of a minimal graph-nonedge pair.
@@
Assume to the contrary that one of the graphs in Figures 20(g) or 20(h) such that $uw_{1}$ is an edge is the expanded $f$-component $H^{-1}$ of some top-level expanded $f$-separating CMS $C^{-1}$ of a minimal graph-nonedge pair $(G,f)$. We will show that either $uw_{1}$ or $w_{2}w_{4}$ is reducing with respect to $(G,f)$, which contradicts the fact that $(G,f)$ is minimal. Observe that $f$ is retained in the minor $[G\cup f]$ obtained by contracting $uw_{1}$, and so $uw_{1}$ is not of Type (3). We attempt to show that $uw_{1}$ is not of Type (1) or (2) by identifying some atom of $[G\cup f]$ contains $[f]$ and has an $[f]$-preserving forbidden minor. Let $G^{\prime}=G\setminus(H^{-1}\setminus C^{-1})$. It suffices to show that some atom of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$ contains $w_{1}w_{4}$ and has a $w_{1}w_{4}$-preserving forbidden minor. By Lemma 33, there exists a nonedge $f^{\prime}$ of $C^{-1}$ such that some atom $J$ of $G^{\prime}\cup f^{\prime}$ contains $f^{\prime}$ and has an $f^{\prime}$-preserving forbidden minor and the pair $(G^{\prime},f^{\prime})$ has no reducing edge. If $f=w_{1}w_{3}$, then the argument is identical to the proof of Lemma 57, so we omit the details.
@@
Next, consider the case where $f^{\prime}=w_{1}w_{4}$. By Lemma 56, $J$ contains $w_{3}$. Hence, since $J$ is an atom, each CMS and atom of $J\cup w_{1}w_{3}$ contains $w_{1}w_{3}$. Combining this with the fact that $J\cup w_{1}w_{3}$ has a $w_{1}w_{4}$-preserving forbidden minor and Lemma 22 shows that some atom of $J\cup w_{1}w_{3}$, and therefore some atom $K$ of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$, contains $w_{1}w_{3}$ and has a $w_{1}w_{4}$-preserving forbidden minor $[K]_{1}$. If $K$ contains $w_{4}$, then the claim is true. Otherwise, we show that $w_{2}w_{4}$ is reducing with respect to $(G,f)$. First, we argue that $K$ has a $w_{1}w_{3}$-preserving forbidden minor. If $[K]_{1}$ preserves $w_{1}w_{3}$, then we are done. Else, since $K$ does not contain $w_{4}$, some CMS $E$ of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$ contained in $K$ is such that all paths between $w_{4}$ and any vertex in $K\setminus E$ contains some vertex in $E$. Also, since $[K]_{1}$ preserves $w_{1}w_{4}$ but not $w_{1}w_{3}$, $E\setminus\{w_{1},w_{3}\}$ contains some vertex $x$ such that $[x]_{1}=[w_{4}]_{1}$. Therefore, there clearly exists a $w_{1}x$-preserving forbidden minor $[K]_{2}$ from which $[K]_{1}$ can be obtained by contracting $[w_{1}w_{3}]_{2}$. The remainder of the proof that $K$ has a $w_{1}w_{3}$-preserving forbidden minor is identical to part of the proof of Lemma 57, so we omit the details.
@@
Finally, observe that $f$ is retained in the minor $[G\cup f]_{4}$ obtained by contracting $w_{2}w_{4}$, and so $w_{2}w_{4}$ is not of Type (3). We complete the proof by showing that $w_{2}w_{4}$ is not of Type (1) or (2). Since $K$ does not contain $w_{4}$, observe that $K$ is an atom of $G\cup w_{1}w_{3}$ and is weakly retained in the minor of $G\cup w_{1}w_{3}$ obtained by contracting $w_{2}w_{4}$. Hence, $K\setminus w_{1}w_{3}$ is some induced subgraph $N$ of $[G^{\prime}]_{4}$ with each vertex $w_{i}$ other than $w_{4}$ as $[w_{i}]_{4}$. Let $G_{1}$ be the subgraph of $[G\cup f]_{4}$ induced by $\{[u]_{4},[w_{1}]_{4},[w_{2}]_{4},[w_{3}]_{4}\}$, $G_{2}=N$, $S=G_{1}\cap G_{2}$, and $e=[w_{1}w_{3}]_{4}$. By construction, $G_{1}\cup G_{2}$ contains $[f]_{4}$. Also, it is easy to see that $G_{1}\cup G_{2}$ has an $[f]_{4}$-preserving forbidden minor. We will apply Lemma 67 to show that $G_{1}\cup G_{2}$ is an atom. Note that $e$ is a nonedge of $G_{1}\cup G_{2}$ and both $G_{1}\cup e$ and $G_{2}\cup e$ are atoms. Furthermore, $G_{1}\setminus A$ is connected for any clique subgraph $A$ of $S$. Lastly, the only clique subgraph of $S$ that could separate $[w_{1}]_{4}$ and $[w_{3}]_{4}$ in $G_{2}$ is $[w_{2}]_{4}$. Since $G_{2}\cup e$ has a forbidden minor and $S$ contains at most three vertices, there exists some vertex in $G_{2}\setminus S$. Additionally, since $G_{2}\cup e$ is an atom, there exist paths that connect this vertex to $[w_{1}]_{4}$ and $[w_{3}]_{4}$ that each contain exactly one vertex in $S$. Therefore, $[w_{2}]_{4}$ does not separate $[w_{1}]_{4}$ and $[w_{3}]_{4}$ in $G_{2}$, and so Lemma 67 shows that $K$ is an atom. Thus, $G_{1}\cup G_{2}$ is contained in some atom of $[G\cup f]_{4}$ that contains $[f]_{4}$ and has an $[f]_{4}$-preserving forbidden minor, and the proof is complete. ∎
@@
First, we show that some atom of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$ contains $w_{1}w_{3}$ and has a $w_{1}w_{3}$-preserving forbidden minor. By Lemma 56, $J$ contains $w_{4}$. Hence, since $J$ is an atom, each CMS and atom of $J\cup w_{1}w_{4}$ contains $w_{1}w_{4}$. Combining this with the fact that $J\cup w_{1}w_{4}$ has a $w_{1}w_{3}$-preserving forbidden minor and Lemma 22 shows that some atom of $J\cup w_{1}w_{4}$, and therefore some atom $K$ of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$, contains $w_{1}w_{4}$ and has a $w_{1}w_{3}$-preserving forbidden minor $[K]_{1}$. If $K$ contains $w_{3}$, then we are done. Otherwise, we show that $w_{2}w_{3}$ is reducing with respect to $(G,f)$.
Observe that $f$ is retained in the minor $[G\cup f]_{1}$ obtained by contracting $w_{2}w_{3}$, and so $w_{2}w_{3}$ is not of Type (3). To see that $w_{2}w_{3}$ is not of Type (1) or (2), note that an argument similar to to part of the proof of Lemma 58 shows that $K$ has a $w_{1}w_{4}$-preserving forbidden minor. Also, since $K$ does not contain $w_{3}$, $K$ is an atom of $G\cup w_{1}w_{4}$ and is weakly retained in the minor of $G\cup w_{1}w_{4}$ obtained by contracting $w_{2}w_{3}$. Hence, $K\setminus w_{1}w_{4}$ is some induced subgraph $N$ of $[G^{\prime}]_{1}$ with each vertex $w_{i}$ other than $w_{3}$ as $[w_{i}]_{1}$. Let $G_{1}=[H^{-1}_{A}$, $G_{2}=N$, $S=G_{1}\cap G_{2}$, and $e=[w_{1}w_{4}]_{1}$. By construction, $G_{1}\cup G_{2}$ contains $[f]_{1}$. Furthermore, it is easy to see that $G_{1}\cup G_{2}$ has an $[f]_{1}$-preserving forbidden minor. Lastly, since $e$ is a nonedge of $G_{1}\cup G_{2}$ and both $G_{1}$ and $G_{2}\cup e$ are atoms, Lemma 66 shows that $G_{1}\cup G_{2}$ is an atom. Therefore, $G_{1}\cup G_{2}$ is contained in some atom of $[G\cup f]_{1}$ that contains $[f]_{1}$ and has an $[f]_{1}$-preserving forbidden minor, and the proof is complete.
Finally, since $K$ contains $w_{3}$, $K\setminus w_{1}w_{3}$ is some induced subgraph $N$ of $[G\cup f]$ with each vertex $w_{i}$ as $[w_{i}]$. Let $G_{1}=[C^{-1}]\cup[u]$, $G_{2}=N$, $S=G_{1}\cap G_{2}$, and $e=[w_{1}w_{3}]$. Observe that $G_{1}\cup G_{2}$ contains $[f]$ and has an $[f]$-preserving forbidden minor. Lastly, since $G_{1}$ and $G_{2}\cup e$ are atoms, Lemma 66 shows that $G_{1}\cup G_{2}$ is an atom. Thus, $G_{1}\cup G_{2}$ is contained in some atom of $[G\cup f]$ that contains $[f]$ and has an $[f]$-preserving forbidden minor, as desired.
@@
By Lemma 56, $J$ contains $w_{2}$ and $w_{3}$. Hence, $J$ is some induced subgraph $G_{2}$ of $[G^{\prime}\cup w_{1}w_{4}]$ with each vertex $w_{i}$ as $[w_{i}]$. Let $G_{1}=[C^{-1}]\cup[u]$, $S=G_{1}\cap G_{2}$, and $e=[w_{1}w_{3}]$. Observe that $G_{1}\cup G_{2}$ contains $[f]$ and has an $[f]$-preserving forbidden minor. Also, since $e$ is a nonedge and both $G_{1}$ and $G_{2}$ are atoms, Lemma 66 shows that $G_{1}\cup G_{2}$ is an atom. Therefore, $G_{1}\cup G_{2}$ is contained in an atom of $[G\cup f]$ that contains $[f]$ and has an $[f]$-preserving forbidden minor. Thus, $xw_{1}$ is reducing, as desired.